If the roots of the equation x2+3x+2 = 0 and x2-x+λ= 0. Are in same ratio then the value of λ ,how to solve

Grade:12

1 Answers

Gurwinder Kaur
13 years ago

x2+3x+2 = 0                           ,,,,              x2-x+λ= 0.

x2+2x+x+2 = 0                       ,,,,               when x=-2 then 4+2+λ=0 ,  λ=-6

x(x+2)+1(x+2)=0                    ,,,,                when x=-1 then 1+1+λ=0  ,  λ=-2

x=-2 , x=-1

plz approve if u like...